8. Given the graph of the function f, use area to compute f(x) dx. y y= f() 3 2 1 2 -1 4 ce
Answer #1

سر f() = Area under the curve from x=0 40x=3 - (Area of the triangle from x=0 to x=1) +( Area of the square from x2 1 to x=3)